# Early Galaxies and Early Dark Energy: A Unified Solution to the Hubble Tension and Puzzles of Massive Bright Galaxies revealed by JWST

### v1

@@ 1 Introduction
The rest-frame UV luminosity functions (UVLFs) determined using JWST-identified galaxies show surprisingly little evolution at the bright end beyond $z\simeq 10$ (e.g. Harikane et al., 2023; Finkelstein et al., 2023b) and differ substantially from the extrapolation of results derived from previous Hubble Space Telescope (HST) observations. Meanwhile, the number densities of UV-bright galaxies at $z\gtrsim 10$ inferred from these observations are in tension with predictions of the majority of the theoretical models (developed before JWST observations, as summarized in e.g. Finkelstein et al., 2023a, b). This includes empirical models (e.g. Tacchella et al., 2013; Mason et al., 2015; Sun & Furlanetto, 2016; Tacchella et al., 2018; Behroozi et al., 2020), semi-analytical galaxy formation models (e.g. Dayal et al., 2014, 2019; Yung et al., 2019; Mauerhofer & Dayal, 2023; Yung et al., 2024b), and cosmological hydrodynamic simulations (e.g. Davé et al., 2019; Vogelsberger et al., 2020; Haslbauer et al., 2022; Kannan et al., 2022, 2023; Wilkins et al., 2023b, a). This discrepancy is largely based on the photometrically-selected galaxy candidates, which could be contaminated by low-redshift interlopers (e.g. Fujimoto et al., 2022; Zavala et al., 2023; Naidu et al., 2022a). However, there has been a good agreement between the photometric and spectroscopic redshifts for the spectroscopically confirmed galaxies so far (e.g. Finkelstein et al., 2023a). The pure spectroscopic constraints of the UVLF (e.g. Harikane et al., 2024) also yield broadly consistent results with the photometric estimates.
@@
Observational signatures of large burstiness of star-formation in high-redshift galaxies have been found (e.g. Ciesla et al., 2023; Cole et al., 2023; Endsley et al., 2023a; Looser et al., 2023; Tacchella et al., 2023a; Dressler et al., 2024; Helton et al., 2024) with a potential qualitative transition around $z\simeq 10$. The major source of this variability could come from the bursty star-formation in low-mass, high-redshift galaxies, often seen in cosmological zoom-in simulations with advanced models for star-formation and feedback in the interstellar medium (ISM; e.g. Hopkins et al., 2018; Faucher-Giguère, 2018; Marinacci et al., 2019; Sun et al., 2023a; Katz et al., 2023). Increased variability at $z\gtrsim 10$ could come from stochastic gas inflow (e.g. Tacchella et al., 2020), rapid star-formation in a feedback-failure regime (e.g. Grudić et al., 2018; Dekel et al., 2023; Menon et al., 2024), in massive star clusters due to Lyman-Werner radiation feedback in metal-poor environment (e.g. Sugimura et al., 2024), or due to highly clustered (in both space and time) feedback from a top-heavy IMF. Enhanced variability in clumpy, high-redshift galaxies could also result indirectly from the limited statistics of star-forming regions and the increased sampling noise.
@@
As a consequence of these changes in cosmological parameters, the halo and thus galaxy abundance are systematically enhanced at high redshifts (Klypin et al., 2021; Boylan-Kolchin, 2023), which has the potential to alleviate the UVLF tension discussed above. The enhanced halo abundance also has implications for another tension raised by JWST regarding potentially overly-massive galaxies at $z\gtrsim 5$ (e.g. Labbé et al., 2023; Akins et al., 2023; Xiao et al., 2023; de Graaff et al., 2024a; Casey et al., 2024). In some cases (e.g. Labbé et al., 2023), the implied cosmic stellar mass density exceeds the total mass budget of baryons in a $\Lambda$CDM universe (Boylan-Kolchin, 2023; Lovell et al., 2023). Despite the debated nature of these sources (e.g. Endsley et al., 2023b; Larson et al., 2022; Kocevski et al., 2023; Desprez et al., 2024; Narayanan et al., 2024; Wang et al., 2024b), EDE offers an alternative way to alleviate this potential cosmological tension.
In this paper, we investigate the potential of EDE as a unified solution to the Hubble tension and the tensions regarding massive, UV-bright galaxies raised by JWST. The paper is organized as follows: In Section 2, we introduce the EDE model and the calculation of dark matter halo mass functions as well as growth rates. In Section 3, we discuss the galaxy formation model and establish a median mapping between the halo mass function and galaxy UVLF. We then describe how we treat UV variability and its halo mass dependence. In Section 4, we present the results and discuss how EDE changes the landscape of UVLFs at $z\gtrsim 10$ and the stellar mass densities at $z\gtrsim 7$. In Section 5, we provide our conclusions.
@@ 2 Cosmological model
The EDE is a scalar field with an axion-like potential $V_{\rm n}(\phi)\sim[1-{\rm cos}(\phi/f)]^{\rm n}$, where $f$ is the decay constant of the field (Poulin et al., 2019). At early times, the field is frozen and acts as a cosmological constant. The field becomes dynamical at $z_{\rm c}$ as the Hubble friction decreases, eventually settling down around the minimum of the potential and starting oscillation. The effective equation-of-state of the field afterward is $w_{\rm a}\simeq(n-1)/(n+1)$. Here following Smith et al. (2022), we consider the $n=3$ case, where the energy density of EDE dilutes faster than radiation. The existence of this additional pre-recombination energy density reduces the physical sound horizon and, in turn, increases the Hubble parameter around the time of photon decoupling. We choose EDE and cosmological parameters as best-fit values in Smith et al. (2022) constrained jointly by ACT, SPT, and Planck results. They are summarized in Table 1 along with the best-fit parameters in the standard $\Lambda$CDM cosmology. The best-fit $H_{0}$ in the EDE cosmology is around $74\,{\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$, in agreement with the local constraints.
@@ 2.1 Halo mass function
The halo mass function is constructed following Press-Schechter-like theories (e.g. Press & Schechter, 1974; Bond et al., 1991; Sheth et al., 2001) as implemented in the Hmf code (Murray et al., 2013; Murray, 2014). The transfer function is calculated using the Code for Anisotropies in the Microwave Background (Camb; Lewis et al. 2000; Howlett et al. 2012) and specifically the axion effective fluid model [^1] We have also experimented with the early quintessence model implemented in Camb and find $\lesssim 0.01$ dex differences between the halo mass function at $z=10$ calculated using the transfer functions from the two EDE model implementations. implemented there. Following Planck Collaboration et al. (2020), in both $\Lambda$CDM and EDE, we model free-streaming neutrinos as two massless and one massive species with $m_{\nu}=0.06\,{\rm eV}$. The effective number of neutrino species $N_{\rm eff}$ is set to $3.046$. We adopt a real-space top-hat filter function for the density field. The definition of halo mass follows the virial criterion in Bryan & Norman (1998). We adopt the halo mass function parametrization of Behroozi et al. (2013) to improve the accuracy at high redshift.
In Figure 1, we show the cumulative number density of haloes above a certain mass threshold as a function of redshift in the EDE and $\Lambda$CDM cosmology. The number densities of haloes are enhanced in the EDE model preferentially at high redshifts and at all mass scales because of the higher values of $n_{\rm s}$, $\omega_{\rm cdm}$ and $\sigma_{8}$ in EDE relative to the Planck $\Lambda$CDM cosmology (see Klypin et al. 2021 and Boylan-Kolchin 2023). By $z=0$, the halo mass functions are indistinguishable.
@@ 3 Empirical galaxy formation model
Figure 3: Star-formation efficiency (SFE) versus halo mass. We show the relation adopted in this work with the gray band, varying $\epsilon$ by ±$\pm 10\%$. We compare it with the SFE in other works (Mason et al., 2015; Tacchella et al., 2018; Behroozi et al., 2019; Harikane et al., 2022; Ferrara et al., 2023b; Li et al., 2023) as labelled. For the Universe Machine (Behroozi et al., 2019) and Ferrara et al. (2023b) predictions, we take their SFR-$M_{\rm halo}$ relations and convert it to SFE assuming our halo accretion rate model. Li et al. (2023) adopted a modified version of Behroozi et al. (2019) with feedback-free starburst (FFB) in massive haloes. For the observational determination in Harikane et al. (2022), we scale it down to account for the different IMF choices.
Figure 4: UV variability $\sigma_{\rm UV}$ versus $M_{\rm halo}$. We show the parametric model adopted in this work, which roughly scales as $M_{\rm halo}^{-1/3}$. Different values of the normalization factor $A$ are preferred at different redshifts beyond $\sim 10$. We compare it to the results from FIRE-2 (Sun et al., 2023a), Sphinx (Katz et al., 2023) as shown in Kravtsov & Belokurov (2024), Serra (Pallottini & Ferrara, 2023) simulations. In both $\Lambda$CDM and EDE, a fixed $\sigma_{\rm UV}$-$M_{\rm halo}$ relation is capable of matching observational results at $z\lesssim 10$. However, at $z\gtrsim 12$, an enhancement of $\sigma_{\rm UV}$ is required and becomes prohibitively high at $z\simeq 16$. In EDE, similar relative enhancement in $\sigma_{\rm UV}$ is inferred at $z\simeq 12-14$ but the overall values of $\sigma_{\rm UV}$ agree reasonably with theoretical predictions. We note that the UVLFs at $z\gtrsim 10$ in the luminosity range probed by observations will not be sensitive to the variability in the massive end, which is indicated by the dashed lines.
@@ 3.1 Star-formation and UV luminosity
This model is a basic representation of our knowledge about galaxy formation before the JWST era. In Figure 3, we compare our model of SFE versus halo mass with other choices in literature, including the Universe Machine predictions at $z\simeq 7-12$, the empirical model in Mason et al. (2015); Tacchella et al. (2018); Ferrara et al. (2023b), the feedback-free starburst (FFB) scenario at $z\simeq 8$ (Li et al., 2023), and the observational constraints in Harikane et al. (2022) based on Halo Occupation Distribution. Our assumed $\epsilon_{\ast}$-$M_{\rm halo}$ relation is a fair representation of the “median” of these models configured before the JWST era. However, we note the substantial uncertainties of the SFE in the low-mass end ($\lesssim 10^{10.5}{\,\rm M_{\odot}}$) and the potential mild increase of SFE at higher redshifts assumed in some of these models. Degeneracy in matching the UVLF does exist between the low-mass end SFE and other factors (e.g. Khimey et al., 2021; Shen et al., 2023; Muñoz et al., 2023), in particular UV variability, which will be discussed in the following sections.
@@ 3.2 UV variability and its dependence on halo mass
where we choose $B=0.34$ which follows Gelli et al. (2024). This has been shown to agree well with the results of cosmological simulations (e.g. Sun et al., 2023a; Katz et al., 2023). $A(z)$ is a free parameter that encapsulates the potential redshift dependence of UV variability. $\sigma_{\rm min}=0.2$ mag is the floor of $\sigma_{\rm UV}$ we introduce better match the low-redshift UVLFs at the bright end (see Section 4.1, although similar agreement can be driven by tuning the massive-end slope $\beta$ of the SFE model). In addition, we assume a maximum value of $\sigma_{\rm max}=2$ mag in low-mass haloes. In Figure 4, we compare this UV variability model at different redshifts with results from various simulations [^2] We adopt a crude estimate of the host halo mass of the two galaxies presented in Pallottini & Ferrara (2023) based on the stellar-to-halo mass relation in Behroozi et al. (2019).  (Sun et al., 2023a; Pallottini & Ferrara, 2023; Katz et al., 2023). We will revisit this figure in Section 4.2 after we examine the UVLFs at different redshifts. In this paper, we have assumed a log-normal distribution of observed UV luminosity. However, in practice, a similar phenomenon can be driven by e.g. incorporating a fraction of starbursts with high SFEs. In Li et al. (2023), a feedback-free starburst (FFB) scenario with $\epsilon^{\rm max}_{\ast}=0.2$ is suggested to explain observational results at $z>10$. Such a scenario can be roughly translated to a feedback-free phase ($\epsilon_{\ast}=1$) with a duty cycle of 0.2 and otherwise the normal phase ($\epsilon_{\ast}\sim 0.02$). The corresponding standard deviation in $\log_{10}{\epsilon_{\ast}}$ is around $0.68$ dex ($1.7$ mag) above the FFB mass-scale $M_{\rm halo}\sim 10^{10}{\,\rm M_{\odot}}$ at $z\gtrsim 12$ (Dekel et al., 2023).

### v2

@@ 1 Introduction
The rest-frame UV luminosity functions (UVLFs) determined using JWST-identified galaxies show surprisingly little evolution at the bright end beyond $z\simeq 10$ (e.g. Harikane et al., 2023; Finkelstein et al., 2023b) and differ substantially from the extrapolation of results derived from previous Hubble Space Telescope (HST) observations. Meanwhile, the number densities of UV-bright galaxies at $z\gtrsim 10$ inferred from these observations are in tension with predictions of the majority of the theoretical models (developed before JWST observations, as summarized in e.g. Finkelstein et al., 2023a, b). This includes empirical models (e.g. Tacchella et al., 2013; Mason et al., 2015; Sun & Furlanetto, 2016; Tacchella et al., 2018; Behroozi et al., 2020), semi-analytical galaxy formation models (e.g. Dayal et al., 2014, 2019; Cowley et al., 2018; Yung et al., 2019; Mauerhofer & Dayal, 2023; Yung et al., 2024b), and cosmological hydrodynamic simulations (e.g. Davé et al., 2019; Vogelsberger et al., 2020; Haslbauer et al., 2022; Kannan et al., 2022, 2023; Wilkins et al., 2023b, a). This discrepancy is largely based on the photometrically-selected galaxy candidates, which could be contaminated by low-redshift interlopers (e.g. Fujimoto et al., 2022; Zavala et al., 2023; Naidu et al., 2022a). However, there has been a good agreement between the photometric and spectroscopic redshifts for the spectroscopically confirmed galaxies so far (e.g. Finkelstein et al., 2023a). The pure spectroscopic constraints of the UVLF (e.g. Harikane et al., 2024b, a) also yield broadly consistent results with the photometric estimates.
@@
Observational signatures of large burstiness of star-formation in high-redshift galaxies have been found (e.g. Ciesla et al., 2023; Cole et al., 2023; Endsley et al., 2023a; Looser et al., 2023; Tacchella et al., 2023a; Dressler et al., 2024; Helton et al., 2024) with a potential qualitative transition around $z\simeq 10$. The major source of this variability could come from the bursty star-formation in low-mass, high-redshift galaxies, often seen in cosmological zoom-in simulations with advanced models for star-formation and feedback in the interstellar medium (ISM; e.g. Hopkins et al., 2018; Faucher-Giguère, 2018; Marinacci et al., 2019; Sun et al., 2023a; Katz et al., 2023). Increased variability at $z\gtrsim 10$ could come from stochastic gas inflow (e.g. Tacchella et al., 2020), rapid star-formation in a feedback-failure regime (e.g. Grudić et al., 2018; Dekel et al., 2023; Menon et al., 2024), in massive star clusters due to Lyman-Werner radiation feedback in metal-poor environment (e.g. Sugimura et al., 2024), or due to highly clustered (in both space and time) feedback from a top-heavy IMF. Enhanced variability in clumpy, high-redshift galaxies could also result indirectly from the limited statistics of star-forming regions and the increased sampling noises.
@@
As a consequence of these changes in cosmological parameters, the halo and thus galaxy abundance are systematically enhanced at high redshifts (e.g. Klypin et al., 2021; Boylan-Kolchin, 2023; Forconi et al., 2024), which has the potential to alleviate the UVLF tension discussed above. The enhanced halo abundance also has implications for another tension raised by JWST regarding potentially overly-massive galaxies at $z\gtrsim 5$ (e.g. Labbé et al., 2023; Akins et al., 2023; Xiao et al., 2023; de Graaff et al., 2024a; Casey et al., 2024). In some cases (e.g. Labbé et al., 2023), the implied cosmic stellar mass density exceeds the total mass budget of baryons in a $\Lambda$CDM universe (Boylan-Kolchin, 2023; Lovell et al., 2023). Despite the debated nature of these sources (e.g. Endsley et al., 2023b; Larson et al., 2022; Kocevski et al., 2023; Desprez et al., 2024; Narayanan et al., 2024; Wang et al., 2024b), EDE offers an alternative way to alleviate this potential cosmological tension.
In this paper, we investigate the potential of EDE as a unified solution to the Hubble tension and the tensions regarding UV-bright and potentially overly massive galaxies raised by JWST. The paper is organized as follows: In Section 2, we introduce the EDE model and the calculation of dark matter halo mass functions as well as growth rates. In Section 3, we discuss the galaxy formation model and establish a median mapping between the halo mass function and galaxy UVLF. We then describe how we treat UV variability and its halo mass dependence. In Section 4, we present the results and discuss how EDE changes the landscape of UVLFs at $z\gtrsim 10$ and the stellar mass densities at $z\gtrsim 7$. In Section 5, we provide discussions and our conclusions.
@@ 2 Cosmological model
The EDE considered in this paper is a scalar field with an axion-like potential $V_{\rm n}(\phi)\sim[1-{\rm cos}(\phi/f)]^{\rm n}$, where $f$ is the decay constant of the field. These ultralight axion-like fields arise generically in string theory (e.g. Svrcek & Witten, 2006; Arvanitaki et al., 2010; Kamionkowski et al., 2014; Marsh, 2016) and have intriguing cosmological implications for dark matter (Marsh, 2016) and EDE (Poulin et al., 2019). The field is frozen at early times and acts as a cosmological constant. The field becomes dynamical at a critical redshift $z_{\rm c}$ as the Hubble friction decreases, eventually settling down around the minimum of the potential and starting oscillation. The effective equation-of-state of the field afterward is $w_{\rm n}\simeq(n-1)/(n+1)$. Here following Smith et al. (2022), we consider the $n=3$ case, where the energy density of EDE dilutes faster than radiation. The existence of this additional pre-recombination energy density increases the Hubble parameter around the time of photon decoupling, and reduces the physical sound horizon if $H_{0}$ is fixed. Therefore, the inferred $H_{0}$ from the same sound horizon measurement on CMB will be increased with EDE. Since the energy density of the EDE field dilutes rapidly, it casts no direct impact on structure formation but its effects are indirectly imprinted through changes in cosmological parameters. We choose EDE and cosmological parameters as the best-fit values in Smith et al. (2022) constrained jointly by ACT, SPT, and Planck results. They are summarized in Table 1 along with the best-fit parameters in the standard $\Lambda$CDM cosmology. The best-fit $H_{0}$ in the EDE cosmology is around $74\,{\rm km}\,{\rm s}^{-1}\,{\rm Mpc}^{-1}$, in agreement with the local constraints.
@@ 2.1 Halo mass function
The halo mass function is constructed following Press-Schechter-like theories (e.g. Press & Schechter, 1974; Bond et al., 1991; Sheth et al., 2001) as implemented in the Hmf code (Murray et al., 2013; Murray, 2014). The transfer function is calculated using the Code for Anisotropies in the Microwave Background (Camb; Lewis et al. 2000; Howlett et al. 2012) and specifically the axion effective fluid model [^1] We have also experimented with the early quintessence model (Smith et al., 2020) implemented in Camb and find $\lesssim 0.01$ dex differences between the halo mass function at $z=10$ calculated using the transfer functions from the two EDE model implementations. (Poulin et al., 2018) implemented there. Following Planck Collaboration et al. (2020), in both $\Lambda$CDM and EDE, we model free-streaming neutrinos as two massless and one massive species with $m_{\nu}=0.06\,{\rm eV}$. The effective number of neutrino species $N_{\rm eff}$ is set to $3.046$. We adopt a real-space top-hat filter function for the density field. The definition of halo mass follows the virial criterion in Bryan & Norman (1998). We adopt the halo mass function parametrization of Behroozi et al. (2013) to improve the accuracy at high redshift.
In Figure 1, we show the cumulative number density of haloes above a certain mass threshold as a function of redshift in the EDE and $\Lambda$CDM cosmology. The number densities of haloes are enhanced in the EDE model preferentially at high redshifts and at all mass scales because of the higher values of $n_{\rm s}$, $\omega_{\rm cdm}$ and $\sigma_{8}$ in EDE relative to the Planck $\Lambda$CDM cosmology (e.g. Klypin et al., 2021; Boylan-Kolchin, 2023; Forconi et al., 2024). By $z=0$, the halo mass functions are indistinguishable.
@@ 3 Empirical galaxy formation model
Figure 3: Star-formation efficiency (SFE) versus halo mass. We show the relation adopted in this work with the gray band, varying $\epsilon$ by ±$\pm 10\%$. We compare it with the SFE in other works (Mason et al., 2015; Tacchella et al., 2018; Behroozi et al., 2019; Harikane et al., 2022; Ferrara et al., 2023b; Li et al., 2023; Ceverino et al., 2024; Feldmann et al., 2024) as labelled. For the Universe Machine (Behroozi et al., 2019) and Ferrara et al. (2023b) predictions, we take their SFR-$M_{\rm halo}$ relations and convert it to SFE assuming our halo accretion rate model. For the FIREbox (Feldmann et al., 2024) predictions, we show the SFE based on the averaged SFR in the 100-Myr window. Li et al. (2023) adopted a modified version of Behroozi et al. (2019) with feedback-free starburst (FFB) in massive haloes. Results from the FirstLight simulations (Ceverino et al., 2024) show signatures of this regime at $z\simeq 12$ (dotted cyan line). For the observational determination in Harikane et al. (2022), we scale it down to account for the different IMF choices.
Figure 4: UV variability $\sigma_{\rm UV}$ versus $M_{\rm halo}$. We show the parametric model adopted in this work, which roughly scales as $M_{\rm halo}^{-1/3}$. Different values of the normalization factor $A$ are preferred at different redshifts beyond $\sim 10$. We compare it to the results from FIRE-2 zoom-in simulations (Sun et al., 2023a), the FIREbox simulation (Feldmann et al., 2024), the Sphinx simulations (Katz et al., 2023) as shown in Kravtsov & Belokurov (2024), and the Serra simulations (Pallottini & Ferrara, 2023). In both $\Lambda$CDM and EDE, a fixed $\sigma_{\rm UV}$-$M_{\rm halo}$ relation is capable of matching observational results at $z\lesssim 10$. However, at $z\gtrsim 12$, an enhancement of $\sigma_{\rm UV}$ is required and becomes prohibitively high at $z\simeq 16$. In EDE, similar relative enhancement in $\sigma_{\rm UV}$ is inferred at $z\simeq 12-14$ but the overall values of $\sigma_{\rm UV}$ agree reasonably with theoretical predictions. We note that the UVLFs at $z\gtrsim 10$ in the luminosity range probed by observations will not be sensitive to the variability in the massive end, which is indicated by the dashed lines.
@@ 3.1 Star-formation and UV luminosity
This model is a basic representation of our knowledge about galaxy formation before the JWST era. In Figure 3, we compare our model of SFE versus halo mass with other choices in literature, including the observational constraints in Harikane et al. (2022) based on Halo Occupation Distribution, the Universe Machine predictions at $z\simeq 7-12$, the empirical models in Mason et al. (2015); Tacchella et al. (2018); Ferrara et al. (2023b), results from the FIREbox simulation (Feldmann et al., 2024), the feedback-free starburst (FFB) scenario at $z\simeq 10$ (Li et al., 2023), and results from the FirstLight simulations (Ceverino et al., 2024, which aligns better with the FFB scenario at $z\gtrsim 10$). Our assumed $\epsilon_{\ast}$-$M_{\rm halo}$ relation is a fair representation of the “median” of the models configured before the JWST era. However, we note the substantial uncertainties of the SFE (e.g. $\epsilon_{\ast}$ can vary between $\sim 1-10\%$ at $M_{\rm halo}\sim 10^{10.5}{\,\rm M_{\odot}}$ among models without entering the FFB regime) and the potential mild increase of SFE at higher redshifts found/assumed in some of these studies (e.g. Behroozi et al., 2019; Ceverino et al., 2024). Degeneracy in matching the UVLF does exist between the low-mass end SFE and other factors (e.g. Khimey et al., 2021; Shen et al., 2023; Muñoz et al., 2023), in particular UV variability, which will be discussed in the following sections.
@@ 3.2 UV variability and its dependence on halo mass
where we choose $B=0.34$ which follows Gelli et al. (2024). This has been shown to agree well with the results of cosmological simulations (e.g. Sun et al., 2023a; Katz et al., 2023). $A(z)$ is a free parameter that encapsulates the potential redshift dependence of UV variability. $\sigma_{\rm min}=0.2$ mag is the floor of $\sigma_{\rm UV}$ we introduce better match the low-redshift UVLFs at the bright end (see Section 4.1, although similar agreement can be driven by tuning the massive-end slope $\beta$ of the SFE model). In addition, we assume a maximum value of $\sigma_{\rm max}=2$ mag in low-mass haloes. In Figure 4, we compare this UV variability model at different redshifts with results from various simulations [^2] We adopt a crude estimate of the host halo mass of the two galaxies presented in Pallottini & Ferrara (2023) based on the stellar-to-halo mass relation in Behroozi et al. (2019).  (Sun et al., 2023a; Pallottini & Ferrara, 2023; Katz et al., 2023; Feldmann et al., 2024). As will be later introduced in Section 4.1 and 4.2, we identify the normalization term $A(z)$ to match the observed UVLFs at different redshifts. We will revisit this figure in Section 4.2 when we discuss the UVLFs at $z\gtrsim 10$.
In this paper, we have assumed a log-normal distribution of observed UV luminosity. However, in practice, a similar phenomenon can be driven by e.g. incorporating a fraction of starbursts with high SFEs. In Li et al. (2023), a feedback-free starburst (FFB) scenario with $\epsilon^{\rm max}_{\ast}=0.2$ is suggested to explain observational results at $z\gtrsim 10$. Such a scenario can be roughly translated to a feedback-free phase ($\epsilon_{\ast}=1$) with a duty cycle of 0.2 and otherwise the normal phase ($\epsilon_{\ast}\sim 0.02$). The corresponding standard deviation in $\log_{10}{\epsilon_{\ast}}$ is around $0.68$ dex ($1.7$ mag) above the FFB mass-scale $M_{\rm halo}\sim 10^{10}{\,\rm M_{\odot}}$ at $z\gtrsim 12$ (Dekel et al., 2023).